OpenStudy (anonymous):

Let f(x) = -4x + 7 and g(x) = 10x - 6, find f(g(x)).

OpenStudy (larseighner):

Work from the inside out. The inside is 10x - 6. Substitute that in the next function outside. f(10x - 6) = -4(10x - 6) + 7 = -40x + 24 + 7 = -40x +31 f(gx) = -40x + 31

OpenStudy (anonymous):

thank you